OpenStudy (anonymous):

3^squareroot -3 3^squareroot15 = 3^squareroot15 True or false?

OpenStudy (anonymous):

\[\sqrt[3]{-3} \sqrt[3]{-5} =\sqrt[3]{15} true or false?\]

OpenStudy (jdoe0001):

\(\bf \huge \sqrt[3]{-3}\sqrt[3]{-5} = \sqrt[3]{15} \ \ ?\)

OpenStudy (anonymous):

True because \(\sqrt[\Large3]{x}\) is an odd-root. It would be false if it were \(\sqrt{x},~\sqrt[\Large4]{x},~\sqrt[\Large6]{x}\), etc.

OpenStudy (anonymous):

Thank you! That's what I thought!